OpenStudy (anonymous):

b. Substitute the result you found from part a. into the equation R = xp to find the revenue equation. Provide your answer in simplified form. This was part a: The final equation would be P=-x+62

OpenStudy (anonymous):

P=-x+62 P-62=-x 62-P=x R=p(62-P)